# Revisiting Meta-evaluation for Grammatical Error Correction

### v1

@@ Abstract
Metrics are the foundation for automatic evaluation in grammatical error correction (GEC), with their evaluation of the metrics (meta-evaluation) relying on their correlation with human judgments. However, conventional meta-evaluations in English GEC encounter several challenges including biases caused by inconsistencies in evaluation granularity, and an outdated setup using classical systems. These problems can lead to misinterpretation of metrics and potentially hinder the applicability of GEC techniques. To address these issues, this paper proposes SEEDA, a new dataset for GEC meta-evaluation. SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems including large language models (LLMs), and two human corrections with different focuses. The results of improved correlations by aligning the granularity in the sentence-level meta-evaluation, suggest that edit-based metrics may have been underestimated in existing studies. Furthermore, correlations of most metrics decrease when changing from classical to neural systems, indicating that traditional metrics are relatively poor at evaluating fluently corrected sentences with many edits.
@@ 1 Introduction
GEC metrics are categorized into edit-based and sentence-based types according to their evaluation granularity, and each has its objectives. The Edit-Based Metric (EBM), such as $M^{2}$ Dahlmeier and Ng (2012) and ERRANT Bryant et al. (2017), focuses on evaluating the quality of the edit itself, whereas the Sentence-Based Metric (SBM), such as GLEU Napoles et al. (2015), evaluates the quality of the entire sentence after correction. Since the system output consists only of sentences without explicit edits, EBM requires the edit extraction from the system output using any method. In addition, these metrics are primarily evaluated based on the correlation with human judgment (i.e., meta-evaluation).
@@
Table 1: Actual data taken from our dataset shows that the results of human evaluation vary depending on the granularity. In edit-based human evaluation, output B was assigned the highest rank, while in sentence-based evaluation, output B received the lowest rank. The results suggest that, even if all edits are considered valid, there are instances where the corrected sentence may lack fluency and naturalness in context when examined as a whole.
@@
Most of the previous meta-evaluations in English GEC have relied on Grundkiewicz et al. (2015)’s dataset with human judgments (henceforth, this dataset is referred to as GJG15). However, existing meta-evaluations based on GJG15 Grundkiewicz et al. (2015); Chollampatt and Ng (2018b); Yoshimura et al. (2020); Gong et al. (2022) have several significant issues. First, the performance of EBM may be underestimated due to biases resulting from inconsistencies in evaluation granularity. As an example of biases, while EBM assigns the lowest score (or the highest score in the sentence-level evaluation) to the uncorrected sentence, sentence-based human evaluation, such as GJG15, assigns scores across the entire range. Furthermore, according to the actual data in Table 1, since human evaluations may yield different results based on granularity, the GEC evaluation suggests a need to separate evaluations for edits and sentences. Second, GJG15 is manually evaluated against the set of classical systems in CoNLL-2014 shared task, such as statistical machine translation approach Junczys-Dowmunt and Grundkiewicz (2014), and classifier-based approach Rozovskaya et al. (2014). Therefore, the gap between the classical systems in GJG15 and the current modern GEC systems based on deep neural networks limits the applicability of meta-evaluation. Third, a single correlation from the current fixed set of systems may not sufficiently capture the performance of metrics, leading to the possibility of drawing incorrect conclusions. For example, Deutsch et al. (2021)’s study on meta-evaluation of summarization revealed that certain metrics can exhibit a spectrum of correlation values, ranging from weak negative to strong positive correlations. Mathur et al. (2020)’s study also showed that outlier systems have a strong influence on correlations in a meta-evaluation of machine translation. Therefore, we are concerned that a similar scenario could occur in the GEC.
@@
To address these issues, we propose SEEDA,  [^1] SEEDA stands for Sentence-based and Edit-based human Evaluation DAtaset for GEC. We have made this dataset publicly available at https://github.com/tmu-nlp/SEEDA. a new dataset to improve the validity of meta-evaluation in English GEC. Specifically, we carefully designed SEEDA to address the first and second issues by performing human evaluations corresponding to two different granularity metrics (i.e., EBM and SBM), covering 12 state-of-the-art system corrections including large language models (LLMs), and two human corrections with different focuses (§3 and §4). Also, through meta-evaluation using SEEDA, we investigate whether EBM, such as $M^{2}$ and ERRANT, is underestimated and demonstrate how the correlation varies between classical systems and neural systems (§6). Furthermore, to address the third issue, we investigate the inadequacy of GEC meta-evaluation based solely on a single correlation by analyzing the presence of outliers and using window analysis (§7). Finally, we discuss best practices and provide recommendations for future researchers to properly meta-evaluate GEC metrics and evaluate their GEC models (§8).
@@ 2 Related work
In the evaluation of GEC, commonly used metrics rely on reference sentences. Some of the most prevalent metrics include $M^{2}$, ERRANT, and GLEU. Both $M^{2}$ and ERRANT calculate an $F_{0.5}$ score by comparing the edits in the corrected sentence to those in the reference. In contrast, GLEU assesses based on the matching of N-grams between the corrected sentence and reference. I-measure evaluates the degree of improvement from the original sentence using the weighted precision of edits. There are also newer metrics like GoToScorer Gotou et al. (2020), which takes into account the difficulty of corrections, and PT-$M^{2}$ Gong et al. (2022), which extends $M^{2}$ (and ERRANT) with pretraining-based metrics. It is worth noting that these reference-based evaluations can lose validity with limited reference coverage.
@@ 3.1 GEC Systems
Figure 1 shows the $M^{2}$ Score ($F_{0.5}$ [^2] In GEC, it is common to use the $F_{0.5}$, where Precision is given twice the importance of Recall Ng et al. (2014); Bryant et al. (2019). This is because, in the practical usage of GEC systems, not correcting is not as detrimental as making incorrect corrections. Additionally, in the context of language acquisition where minimizing incorrect feedback is desirable, this weighting is reasonable Nagata and Nakatani (2010). and word edit rate for classical systems in GJG15, neural systems in SEEDA, and human sentences. Comparing these systems, neural systems in SEEDA show a higher number of edits and demonstrate better correction performance from the perspective of $M^{2}$. This performance comparison utilizes the most common GEC evaluation method, reproducing results reported in existing studies. On the other hand, this performance comparison contains intuitive contradictions, such as the lower performance of human-corrected sentences and LLMs. Therefore, we investigate and report how the modern system comparison deviates from human judgments (§4.2). Note that few-shot learning such as GPT-3.5 is known to be not grounding to target sentences as compared to finetuned models and may produce fluent but lengthy correction sentences that do not preserve the meaning of the source Maynez et al. (2023).
@@ 3.2 Annotation scheme
Figure 2 shows an overview of the annotation flow and an example of edit-based human evaluation. In Step 1, the detection of errors in the source and checking for edits in the output are performed. During the initial error detection, annotators refer to 25 error categories by Bryant et al. (2017) to identify error locations in the source, enabling them to label errors at the minimal unit level. In the subsequent Edit checking, annotators perform a binary decision to determine whether they would like to apply the edits in the output to improve the source or not. To reduce annotation costs, ERRANT is used for extracting edits. When there are conflicting edits (e.g., subject-verb agreement error), the one that aligns with the context is deemed effective, while the other is considered ineffective. Furthermore, for edits that depend on each other (e.g., [law’s→ ] and [ →of the law] in Figure 2), each is assigned an independent label, but they are deemed effective only if all dependent edits are present. In Step 2, the annotator performs a binary decision to determine whether each edit in the output effectively corrects the errors found in Step 1. Finally, we compute the $F_{0.5}$ based on Precision and Recall for each corrected sentence and subsequently rank the set of corrected sentences accordingly. The supplementary information about the annotation is provided in Appendix A.
@@
 # Score Range System 1 0.679 1 REF-F 2 0.583 2 GPT-3.5 3 0.173 3 TransGEC 4 0.097 4-6 T5 0.078 4-7 REF-M 0.067 4-7 Riken-Tohoku 0.064 4-7 BERT-fuse 5 -0.076 8-11 UEDIN-MS -0.084 8-11 PIE -0.092 8-11 GECToR-BERT -0.097 8-11 LM-Critic 6 -0.154 12-12 GECToR-ens 7 -0.211 13-14 TemplateGEC -0.231 13-14 BART 8 -0.797 15 INPUT
(a) Sentence-based evaluation in GJG15
@@ 4.1 Dataset statistics
Figure 3: Scatter plots of the human score and the metric score. “Base” indicates the 12 systems excluding uncorrected sentences (INPUT) and fluent sentences (REF-F, GPT-3.5). Each line represents a regression line, and the shaded area indicates the size of the confidence interval for the estimated regression, obtained using bootstrap. Comparing the red and green regression lines to the blue regression line allows us to observe the degree of influence of each outlier on the distribution trend. For example, the leftward tilt of the red regression lines for $M^{2}$, PT-$M^{2}$, ERRANT, and GLEU indicates a negative impact from fluent sentences as outliers.
@@ 5 Baseline metrics
We target 11 GEC metrics for meta-evaluation, including EBM (§ 5.1) and SBM (§ 5.2).
@@ 5.1 Edit-based metrics
It is similar to $M^{2}$ but differs in that it uses linguistically enhanced Damerau-Levenshtein alignment for extracting edits. It is characterized by its ability to calculate $F_{0.5}$ scores for each error type.
@@
It is a variant of ERRANT that computes sentence-level $F_{0.5}$ scores.
@@
It calculates the $F_{0.5}$ score while considering the difficulty of correction. The difficulty is calculated based on the number of systems that were able to correct the error.
@@ 6.1 System-level meta-evaluation
For our system-level meta-evaluation, in Table 2, we report correlation using system scores obtained from human rankings. Metrics such as $M^{2}$, PT-$M^{2}$, ERRANT, GoToScorer, and GLEU can calculate system scores, while other metrics use the average of sentence-level scores as the system score. We use Pearson correlation (r) and Spearman rank correlation ($\rho$) to measure the closeness between the metric and human evaluation.
@@
According to the system-level meta-evaluation results in Table 5, it is evident that aligning the granularity between the metrics and human evaluation improves the correlation for EBM in SEEDA-E, while the correlation for SBM in SEEDA-S tends to decrease. One reason for the inconsistent results even when the granularity is aligned is that system-level human evaluations exhibit relatively small variations across different evaluation granularities.
We discovered that as we move from classical systems to neural systems, correlations for all metrics—except GoToScorer and GLEU—decrease through a comparison between GJG15 and SEEDA-S. This result suggests that the majority of current metrics cannot adequately evaluate the more extensively edited and fluent corrections produced by neural systems, in contrast to those generated by classical systems. In the meta-evaluation results of GJG15, comparing it with existing studies Grundkiewicz et al. (2015); Choshen and Abend (2018a) that is unfeasible, as the exclusion of INPUT has been implemented to alleviate scoring bias between EBM and sentence-based human evaluation.
Figure 4: Variation of correlation when different systems are considered using window analysis. The x-axis represents the human ranking of the 12 systems excluding outliers. “n” denotes the number of systems considered, with solid lines representing four systems and dashed lines representing eight systems. For example, for n$=$4, a point with x$=$5 corresponds to a human evaluation using systems ranked 2 to 5. The red line represents Pearson (r) and the blue line represents Kendall ($\tau$). The correlation of the main metrics in GEC ($M^{2}$, ERRANT, GLEU) shows significant variability, while pretraining-based metrics (SOME, IMPARA) exhibit relatively stable correlations.
@@ 6.2 Sentence-level meta-evaluation
When we compared GJG15 and SEEDA-S, we observed a decrease in correlations for most metrics, especially in EBM, similar to the system-level results. Consistently high correlations were found for SOME and IMPARA, indicating the effectiveness of fine-tuned BERT.
@@ 7.1 Influence of outliers
The improved correlations in $M^{2}$ (+Min) and GLEU (+Min) when REF-F and GPT-3.5 were added indicate that the fluency correction may no longer be an outlier for commonly used metrics if the low coverage of reference-based evaluation is mitigated. To address the issue of reference coverage, an approach similar to Choshen and Abend (2018a), which involves splitting and combining edits for each reference, could potentially enhance the effective utilization of references. However, the result that fluency edit references were useful only for GLEU suggests that fluent edit references may be effective on an N-gram basis, but not on an edit extraction basis. As one of the reasons, we can consider the difficulties and complexities in edit extraction for fluent sentences in EBM, as well as the inability to address the low coverage of three fluent references. The fact that SOME and IMPARA maintained high correlations suggests that there is a wide scope of corrections that can be properly evaluated in GEC.
@@ 8.1 Towards valid meta-evaluation in GEC
We recommend that meta-evaluation be conducted at each evaluation granularity in GEC. Specifically, EBM should use SEEDA-E, and SBM should use SEEDA-S. The meta-evaluation using SEEDA should use the 12 systems as a baseline, excluding outliers, and add REF-F and GPT-3.5 if you want to find out how well the metrics can evaluate fluent corrections. This allows meta-evaluation for the modern neural system without the bias of the granularity. Additionally, conducting experiments with various methodologies is crucial to validate the characteristics of metrics. Therefore, experiments using GMEG-Data for domain-specific meta-evaluation of SBM and meta-evaluation by MAEGE, irrespective of granularity, should be considered if resources permit.

### v2

@@ Abstract
Metrics are the foundation for automatic evaluation in grammatical error correction (GEC), with their evaluation of the metrics (meta-evaluation) relying on their correlation with human judgments. However, conventional meta-evaluations in English GEC encounter several challenges including biases caused by inconsistencies in evaluation granularity, and an outdated setup using classical systems. These problems can lead to misinterpretation of metrics and potentially hinder the applicability of GEC techniques. To address these issues, this paper proposes SEEDA, a new dataset for GEC meta-evaluation. SEEDA consists of corrections with human ratings along two different granularities: edit-based and sentence-based, covering 12 state-of-the-art systems including large language models (LLMs), and two human corrections with different focuses. The results of improved correlations by aligning the granularity in the sentence-level meta-evaluation suggest that edit-based metrics may have been underestimated in existing studies. Furthermore, correlations of most metrics decrease when changing from classical to neural systems, indicating that traditional metrics are relatively poor at evaluating fluently corrected sentences with many edits.
@@ 1 Introduction
GEC metrics are categorized into edit-based and sentence-based types according to their evaluation granularity, and each has its objectives. Edit-Based Metrics (EBMs), such as $M^{2}$ Dahlmeier and Ng (2012) and ERRANT Bryant et al. (2017), focus on evaluating the quality of the edit itself, whereas Sentence-Based Metrics (SBMs), such as GLEU Napoles et al. (2015), evaluate the quality of the entire sentence after correction. Since the system output consists only of sentences without explicit edits, EBMs require the edit extraction from the system output using any method. In addition, these metrics are primarily evaluated based on the correlation with human judgment (i.e., meta-evaluation).
@@
Table 1: Actual data taken from our dataset shows that the results of human evaluation vary depending on the granularity. In edit-based evaluation, output B was assigned the highest rank (tied with output A), while in sentence-based evaluation, output B received the lowest rank. The results suggest that, even if all edits are considered valid, there are instances where the corrected sentence may lack fluency and naturalness in context.
@@
Most of the previous meta-evaluations in English GEC have relied on Grundkiewicz et al. (2015)’s dataset with human judgments (henceforth, this dataset is referred to as GJG15). However, existing meta-evaluations based on GJG15 Grundkiewicz et al. (2015); Chollampatt and Ng (2018b); Yoshimura et al. (2020); Gong et al. (2022) have several significant issues. First, the performance of EBMs may be underestimated due to biases resulting from inconsistencies in evaluation granularity. As an example of biases, while EBMs assign the lowest score (or the highest score in the sentence-level evaluation) to the uncorrected sentence, sentence-based human evaluation, such as GJG15, assigns scores across the entire range. Furthermore, according to the actual data in Table 1, since human evaluations may yield different results based on granularity, the GEC evaluation suggests a need to separate evaluations for edits and sentences. Second, GJG15 is manually evaluated against the set of classical systems in CoNLL-2014 shared task, such as statistical machine translation approach Junczys-Dowmunt and Grundkiewicz (2014), and classifier-based approach Rozovskaya et al. (2014). Therefore, the gap between the classical systems in GJG15 and the current modern GEC systems based on deep neural networks limits the applicability of meta-evaluation. Third, a single correlation from the current fixed set of systems may not sufficiently capture the performance of metrics, leading to the possibility of drawing incorrect conclusions. For example, Deutsch et al. (2021)’s study on meta-evaluation of summarization revealed that certain metrics can exhibit a spectrum of correlation values, ranging from weak negative to strong positive correlations. Mathur et al. (2020)’s study also showed that outlier systems have a strong influence on correlations in a meta-evaluation of machine translation. Therefore, we are concerned that a similar scenario could occur in the GEC.
@@
To address these issues, we propose SEEDA [^1] SEEDA stands for Sentence-based and Edit-based human Evaluation DAtaset for GEC. We have made this dataset publicly available at https://github.com/tmu-nlp/SEEDA. a new dataset to improve the validity of meta-evaluation in English GEC. Specifically, we carefully designed SEEDA to address the first and second issues by performing human evaluations corresponding to two different granularity metrics (i.e., EBMs and SBMs), covering 12 state-of-the-art system corrections including large language models (LLMs), and two human corrections with different focuses (§3 and §4). Also, through meta-evaluation using SEEDA, we investigate whether EBMs, such as $M^{2}$ and ERRANT, are underestimated and demonstrate how the correlation varies between classical systems and neural systems (§6). Furthermore, to address the third issue, we investigate the inadequacy of GEC meta-evaluation based solely on a single correlation by analyzing the presence of outliers and using window analysis (§7). Finally, we discuss best practices and provide recommendations for future researchers to properly meta-evaluate GEC metrics and evaluate their GEC models (§8).
@@ 2 Related work
In the evaluation of GEC, commonly used metrics rely on reference sentences. Some of the most prevalent metrics include $M^{2}$, ERRANT, and GLEU. Both $M^{2}$ and ERRANT calculate $F_{0.5}$ score by comparing the edits in the corrected sentence to those in the reference. In contrast, GLEU assesses based on the matching of N-grams between the corrected sentence and reference. I-measure evaluates the degree of improvement from the original sentence using the weighted precision of edits. There are also newer metrics like GoToScorer Gotou et al. (2020), which takes into account the difficulty of corrections, and PT-$M^{2}$ Gong et al. (2022), which extends $M^{2}$ (and ERRANT) with pretraining-based metrics. It is worth noting that these reference-based evaluations can lose validity with limited reference coverage.
@@ 3.1 GEC Systems
Figure 1 shows the $M^{2}$ Score ($F_{0.5}$ [^2] In GEC, it is common to use $F_{0.5}$, where Precision is given twice the importance of Recall Ng et al. (2014); Bryant et al. (2019). This is because, in the practical usage of GEC systems, not correcting is not as detrimental as making incorrect corrections. Additionally, in the context of language acquisition where minimizing incorrect feedback is desirable, this weighting is reasonable Nagata and Nakatani (2010). and word edit rate for classical systems in GJG15, neural systems in SEEDA, and human sentences. Comparing these systems, neural systems in SEEDA show a higher number of edits and demonstrate better correction performance from the perspective of $M^{2}$. This performance comparison utilizes the most common GEC evaluation method, reproducing results reported in existing studies. On the other hand, this performance comparison contains intuitive contradictions, such as the lower performance of human-corrected sentences and LLMs. Therefore, we investigate and report how the modern system comparison deviates from human judgments (§4.2). Note that few-shot learning such as GPT-3.5 is known to be not grounding to target sentences as compared to finetuned models and may produce fluent but lengthy correction sentences that do not preserve the meaning of the source Maynez et al. (2023).
@@ 3.2 Annotation scheme
Figure 2 shows an overview of the annotation flow and an example of edit-based human evaluation. In Step 1, the detection of errors in the source and checking for edits in the output are performed. During the initial error detection, annotators refer to 25 error categories by Bryant et al. (2017) to identify error locations in the source, enabling them to label errors at the minimal unit level. In the subsequent Edit checking, annotators perform a binary decision to determine whether they would like to apply the edits in the output to improve the source or not. To reduce annotation costs, ERRANT is used for extracting edits. When there are conflicting edits (e.g., subject-verb agreement error), the one that aligns with the context is deemed effective, while the other is considered ineffective. Furthermore, for edits that depend on each other (e.g., [law’s→ ] and [ →of the law] in Figure 2), each is assigned an independent label, but they are deemed effective only if all dependent edits are present. In Step 2, the annotator performs a binary decision to determine whether each edit in the output effectively corrects the errors found in Step 1. Finally, we compute $F_{0.5}$ based on Precision and Recall  [^3] Note that Precision and Recall are computed at different levels of granularity. for each corrected sentence and subsequently rank the set of corrected sentences accordingly. The supplementary information about the annotation is provided in Appendix A.
@@
 # Score Range System
1 0.679 1 REF-F
2 0.583 2 GPT-3.5
3 0.173 3 TransGEC
4 0.097 4-6 T5
0.078 4-7 REF-M
0.067 4-7 Riken-Tohoku
0.064 4-7 BERT-fuse
5 -0.076 8-11 UEDIN-MS
-0.084 8-11 PIE
-0.092 8-11 GECToR-BERT
-0.097 8-11 LM-Critic
6 -0.154 12-12 GECToR-ens
7 -0.211 13-14 TemplateGEC
-0.231 13-14 BART
8 -0.797 15 INPUT
@@ 4.1 Dataset statistics
Figure 3: Scatter plots of the human score and the metric score. “Base” indicates the 12 systems excluding uncorrected sentences (INPUT) and fluent sentences (REF-F, GPT-3.5). Each line represents a regression line, and the shaded area indicates the size of the confidence interval for the estimated regression, obtained using bootstrap. Comparing the orange and blue regression lines to the gray regression line allows us to observe the degree of influence of each outlier on the distribution trend. For example, the leftward tilt of the orange regression lines for $M^{2}$, PT-$M^{2}$, ERRANT, and GLEU indicates a negative impact from fluent sentences as outliers.
@@ 5 Baseline metrics
We target 11 GEC metrics for meta-evaluation, including EBMs (§ 5.1) and SBMs (§ 5.2).
@@ 5.1 Edit-based metrics
It is similar to $M^{2}$ but differs in that it uses linguistically enhanced Damerau-Levenshtein alignment for extracting edits. It is characterized by its ability to calculate $F_{0.5}$ score for each error type.
@@
It is a variant of ERRANT that computes sentence-level $F_{0.5}$ score.
@@
It calculates $F_{0.5}$ score while considering the difficulty of correction. The difficulty is calculated based on the number of systems that were able to correct the error.
@@ 6.1 System-level meta-evaluation
For our system-level meta-evaluation, we report correlation using system scores obtained from human rankings. Metrics such as $M^{2}$, PT-$M^{2}$, ERRANT, GoToScorer, and GLEU can calculate system scores, while other metrics use the average of sentence-level scores as the system score. We use Pearson correlation (r) and Spearman rank correlation ($\rho$) to measure the closeness between the metric and human evaluation.
@@
According to the system-level meta-evaluation results in Table 5, it is evident that aligning the granularity between the metrics and human evaluation improves the correlation for EBMs in SEEDA-E, while the correlation for SBMs in SEEDA-S tends to decrease. One reason for the inconsistent results even when the granularity is aligned is that system-level human evaluations exhibit relatively small variations across different evaluation granularities.
We discovered that as we move from classical systems to neural systems, correlations for all metrics—except GoToScorer and GLEU—decrease through a comparison between GJG15 and SEEDA-S. This result suggests that the majority of current metrics cannot adequately evaluate the more extensively edited and fluent corrections produced by neural systems, in contrast to those generated by classical systems. In the meta-evaluation results of GJG15, comparing it with existing studies Grundkiewicz et al. (2015); Choshen and Abend (2018a) is unfeasible, as the exclusion of INPUT has been implemented to alleviate scoring bias between EBMs and sentence-based human evaluation.
Figure 4: Variation of correlation when different systems are considered using window analysis. The x-axis represents the human ranking of the 12 systems excluding outliers. “n” denotes the number of systems considered, with solid lines representing four systems and dashed lines representing eight systems. For example, for n$=$4, a point with x$=$5 corresponds to a human evaluation using systems ranked 2 to 5. The orange line represents Pearson (r) and the blue line represents Kendall ($\tau$). The correlation of the main metrics ($M^{2}$, ERRANT, GLEU) shows significant variability, while pretraining-based metrics (SOME, IMPARA) exhibit relatively stable correlations.
@@ 6.2 Sentence-level meta-evaluation
When we compared GJG15 and SEEDA-S, we observed a decrease in correlations for most metrics, especially in EBMs, similar to the system-level results. Consistently high correlations were found for SOME and IMPARA, indicating the effectiveness of fine-tuned BERT.
@@ 7.1 Influence of outliers
The improved correlations in $M^{2}$ (+Min) and GLEU (+Min), when REF-F and GPT-3.5 were added, indicate that the fluency correction may no longer be an outlier for commonly used metrics if the low coverage of reference-based evaluation is mitigated. To address the issue of reference coverage, an approach similar to Choshen and Abend (2018a), which involves splitting and combining edits for each reference, could potentially enhance the effective utilization of references. However, the result that fluency edit references were useful only for GLEU suggests that fluent edit references may be effective on an N-gram basis, but not on an edit extraction basis. As one of the reasons, we can consider the difficulties and complexities in edit extraction for fluent sentences in EBMs, as well as the inability to address the low coverage of three fluent references.
@@ 8.1 Towards valid meta-evaluation in GEC
We recommend that meta-evaluation be conducted at each evaluation granularity in GEC. Specifically, EBMs should use SEEDA-E, and SBMs should use SEEDA-S. The meta-evaluation using SEEDA should use the 12 systems as a baseline, excluding outliers, and add REF-F and GPT-3.5 if you want to find out how well the metrics can evaluate fluent corrections. This allows meta-evaluation for the modern neural system without the bias of the granularity. Additionally, conducting experiments with various methodologies is crucial to validate the characteristics of metrics. Therefore, experiments using GMEG-Data for domain-specific meta-evaluation of SBMs and meta-evaluation by MAEGE, irrespective of granularity, should be considered if resources permit.